A current of 1000 MA passes through a conductor for 2minute. Calculate quantity of electricity transfered
1
Expert's answer
2020-11-25T10:45:23-0500

Solution

Current I=109A

Time t=120 sec

Current "I = \frac{Q}{t}"

Transffered charge

"Q =It=10^9\times 120= 12\times10^{10}"C